OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-5, 10) and (-4, 3). Please show all of your work.

OpenStudy (mysesshou):

so , we start with this format http://en.wikipedia.org/wiki/Linear_equation#Two-point_form

OpenStudy (mysesshou):

after finding the slope

OpenStudy (mysesshou):

passes through (-5, 10) and (-4, 3). to find slope: the change of Y over the change in X \[m= \frac{10-3}{(-5)-(-4)}= \frac{7}{(-5)+4}= \frac{7}{-1} \ -7\] ok?

OpenStudy (mysesshou):

that ok so far?

OpenStudy (anonymous):

yes

OpenStudy (mysesshou):

Then, use the format of the equation from wiki \[(y - y1) = m* (x - x1)\] where x1 and y1 are one of the points, (x1,y1), let's pick (-5,10), but it doesn't really matter which. So, substitute in the y1, x1, m into the equation: \[(y - 10) = (-7)* (x - (-5))\] then...

OpenStudy (mysesshou):

simplify: y−10 = (−7)∗(x−(−5)) the subtraction of a negative number changes it to addition y−10 = (−7)∗(x+5) multiply through the -7 y−10 = −7x-35 add 10 to both sides y = −7x-25 so y = −7x-25 is the slope intercept form -7 is the slope -25 is the intercept y = (−7)*x+(-25) y = (m) *x + (b)

OpenStudy (anonymous):

i am writing down the steps

OpenStudy (mysesshou):

ok. let me check that

OpenStudy (mysesshou):

ok, works

OpenStudy (anonymous):

so my final equation would be....

OpenStudy (mysesshou):

haven't seen it yet. unless i missed it. :)

OpenStudy (mysesshou):

ok, write when you go post it. :)

OpenStudy (anonymous):

ok i didnt get it all the way as im writing it it looks weird lol

OpenStudy (mysesshou):

ok, let me know what you need. not sure how to explain more through typing

OpenStudy (anonymous):

okay lets do the last one then we are done woooo hoooo

OpenStudy (mysesshou):

is it on th left?

OpenStudy (anonymous):

yes